Hero:

How long would it take to read 37 thousand pages through Text To Speech?

Ultrilliam:

Welp, let's calculate this out. *cracks knuckles*

SmokeyBrown:

Wait, Windows 10 has an option to change the speed of the text-to-speech voice. And many different voices to choose from. Are we assuming default settings?

Hero:

The average WPM for TTS is 200.

Hero:

Hopefully Ultrilliam factored that into his calculations.

Ultrilliam:

I'd say the average text to speech would take about 4 minutes and 30 seconds to read a page. So lets convert it to seconds for simplicity 60*4=240+30=270 Now let's multiply that by 37000, 270*37000=9990000 Let's divide that by 60 to get that in minutes 9990000/60=166500 and now by 60 again to get hours 166500/60=2775 And now by 24 to get days 115.625 days. This however does not factor in sleeping, if we were to factor in that the average person sleeps for 8 hours a day. 2775*8=22200 We are getting closer, but the average person also eats for about 1 hour and 30 minutes a day, let's convert both numbers to minutes and multiply them 22200*60=1332000*90=119880000/60/24/365=200+ years I done took a wrong turn here somewhere...

Ultrilliam:

Probably in adding hours you wouldn't be doing it I done goofed. Hm

SmokeyBrown:

I think you could simplify the calculations by assuming the number of hours per day you'd be running the software and dividing by that number instead of 24 in the hours-to-days conversion. So, 16 hours a day (8 hours of sleeping) would just be 2775/16, or 173.43

Ultrilliam:

^ah, much more elegant

SmokeyBrown:

If we take away 1.5 hours for eating, we get 2775/14.5, or 191.37 days

Ultrilliam:

Let's also take 30 minutes away for showering =P 2775/14 or 198.21 days

SmokeyBrown:

I guess Hero deserves some input here. How many hours would you like us to assume per day, Hero? Although, you could probably crunch the numbers yourself at this point :p

Hero:

It wasn't as difficult as you made it out to be. 128 days PLUS you factor in to your calculations the time spent sleeping. 4.5 pages per minute. For simplicity, I'm rounding that up to 5 pages per minute. Therefore it will take 37,000/5 = 7400 minutes to read the whole thing plus the time it takes to sleep. 7400 minutes divided by 1400 minutes per day comes out to about 5.13 days not including the time spent sleeping.

Hero:

I know that sounds shocking but that's what the numbers say so far. Now to add the time spent sleeping to the total amount of days ...

Hero:

Yeah, it took me so long because I had to make sure my numbers were correct. They seemed off at first.

Ultrilliam:

4.5 PAGES per MINUTE? I think that was your mess up?

Ultrilliam:

5 minutes per page seems more accurate

Hero:

Whoops. Oh well.

Hero:

It's getting late for me.

Hero:

Bottom line is, no one is going to attempt to read that many pages in a year. Let alone an entire lifetime. It was fun to try calculating this though.